# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@ 1.5 Positive Traces
    If $H(z)$ is a continuous complex-valued function on $\mathbb{R}$ with finitely many zeros and at most polynomial growth at infinity, then the set $\{H(z)S(z):S(z)\in\mathbb{C}[z]\}$ is dense in the space $L^{p}(\mathbb{R},w)$.
@@
    The closure of the set $\{S(z)\bar{S}(z):S(z)\in\mathbb{C}[z]\}$ in $L^{p}(\mathbb{R},w)$ is the subset of almost everywhere nonnegative functions.
We begin by showing that the linear functional $\Phi_{0}$ of Theorem 1.4.3 must be identically zero.
Proposition 1.5.2.
Any positive trace $T$ on $\mathcal{O}_{c}$ of the form given by Theorem 1.4.3 satisfies $\Phi_{0}(R)=0$ for all polynomials $R$.
@@
Since $T$ is a positive trace, we must have $T(a\rho(a))>0$ for all $a$ in $\mathcal{O}_{c}$. In particular, we must have $T(R(k)\bar{R}(k)e)>0$ for all even polynomials $R$. By Condition (0) of Theorem 1.3.2, this trace value is equal to $\frac{1}{2}T(R(k)\bar{R}(k)e_{0})+\frac{1}{2}R(0)\bar{R}(0)T(e_{0}h)$.
Suppose, for the sake of contradiction, that $\Phi_{0}$ is not identically zero. Then there exists a polynomial $S_{0}$ such that $S_{0}(0)=0$ and $\Phi_{0}(S_{0})$ is negative. By Remark 1.4.4, we may take $S_{0}$ to be even. Also, let
@@
where the product is over singularities $z_{0}$ of $w_{0}(z)$ between $i\mathbb{R}$ and the path $C$. so that $H(0)=0$ and $H(\alpha_{i}-i-\frac{1}{2})=0$ for all $i$. Thus for any polynomial $U(z)$, we have $\Phi_{0}((S_{0}-HU)(\bar{S}_{0}-\bar{H}\bar{U}))=\Phi_{0}(S_{0})$ and $S_{0}(0)-H(0)U(0)=0$. Also, by Remark 1.4.4, the polynomial $H(z)$ is even.
By part (1) of Lemma 1.5.1, there exists a sequence of polynomials $S_{i}$ such that $HS_{i}$ tends to $S_{0}$ in the space $L^{2}(i\mathbb{R},|w_{0}|)$. Because $w_{0}(z)$ is even, the sequence $H(z)\cdot\frac{S_{i}(z)+S_{i}(-z)}{2}$ also tends to $S_{0}$ in $L^{2}(i\mathbb{R},|w_{0}|)$. Thus the integral
$\int_{i\mathbb{R}}\left(S_{0}(z)-H(z)\cdot\frac{S_{i}(z)+S_{i}(-z)}{2}\right)\left(\bar{S}_{0}(z)-\bar{H}(z)\cdot\frac{\bar{S}_{i}(z)+\bar{S}_{i}(-z)}{2}\right)w_{0}(z)\mathrm{d}z$
@@
$T(R(k)\bar{R}(k)e)$ $=\tfrac{1}{2}T(R(k)\bar{R}(k)e_{0})+\tfrac{1}{2}R(0)\bar{R}(0)T(e_{0}h)$
$=\tfrac{1}{2}T(R(k)\bar{R}(k)e_{0})$
$=\int_{i\mathbb{R}}R(z)\bar{R}(z)w_{0}(z)\mathrm{d}z+\Phi_{0}(R),$
@@
$T(a\rho(a))$ $=T\left(R(k)(x^{n}+y^{n})e\bar{R}(k)(x^{n}+y^{n})e\right)$
$=T\left(R(k)(x^{n}+y^{n})\bar{R}(k)(x^{n}+y^{n})e\right)$
$=T\left(R(k)(x^{n}+y^{n})\left(x^{n}\bar{R}(k-n)+y^{n}\bar{R}(k+n)\right)e\right),$
and distributing and using the commutativity relations with $k$ listed in Lemma 1.2.1, we obtain
$T(a\rho(a))$ $=T\left(R(k)\left(x^{2n}\bar{R}(k-n)+x^{n}y^{n}\bar{R}(k+n)+y^{2n}\bar{R}(k+n)+y^{n}x^{n}\bar{R}(k-n)\right)e\right)$
$=T\left(x^{2n}R(k-2n)\bar{R}(k-n)e\right)+T\left(x^{n}y^{n}R(k)\bar{R}(k+n)e\right)$
$\phantom{=\ }+T\left(y^{2n}R(k+2n)\bar{R}(k+n)e\right)+T\left(y^{n}x^{n}R(k)\bar{R}(k-n)e\right).$
@@
$T\left(a\rho(a)\right)$ $=\tfrac{1}{2}T\left(x^{2n}R(k-2n)\bar{R}(k-n)e_{0}h\right)+\tfrac{1}{2}T\left(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}h\right)$
$\phantom{=\ }+\tfrac{1}{2}T\left(y^{2n}R(k+2n)\bar{R}(k+n)e_{0}h\right)+\tfrac{1}{2}T\left(y^{n}x^{n}R(k)\bar{R}(k-n)e_{0}h\right)$
$\phantom{=\ }+\tfrac{1}{2}T\left(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}\right)+\tfrac{1}{2}T\left(y^{n}x^{n}R(k)\bar{R}(k-n)e_{0}\right).$
Using Conditions (0) and (0) of Theorem 1.3.2 and rearranging terms, we obtain
$T(a\rho(a))$ $=\tfrac{1}{2}T\left(x^{n}y^{n}R(k-n)\bar{R}(k)e_{0}h\right)+\tfrac{1}{2}T\left(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}h\right)$
$\phantom{=\ }+\tfrac{1}{2}T\left(y^{n}x^{n}R(k+n)\bar{R}(k)e_{0}h\right)+\tfrac{1}{2}T\left(y^{n}x^{n}R(k)\bar{R}(k-n)e_{0}h\right)$
$\phantom{=\ }+\tfrac{1}{2}T\left(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}\right)+\tfrac{1}{2}T\left(y^{n}x^{n}R(k)\bar{R}(k-n)e_{0}\right).$
@@
$T\left(a\rho(a)\right)$ $=T\left(x^{n}y^{n}e_{0}h\right)(R(n)\bar{R}(0)+R(0)\bar{R}(n))+T\left(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}\right)$
$=T(e_{0}h)(R(n)\bar{R}(0)+R(0)\bar{R}(n))\prod\limits_{q=0}^{n-1}(c_{0}q+\beta_{q})$
$\phantom{=\ }+\int_{i\mathbb{R}}R(z)\bar{R}(z+n)\left(\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})\right)w_{0}(z)|\mathrm{d}z|.$ (1.2)
@@
The product $\prod_{q=0}^{n-1}(c_{0}q+\beta_{q})$ in the first term is nonzero by assumption; indeed, by Lemma 1.2.5, $c_{0}q+\beta_{q}=c_{0}(q+\frac{1}{2}-\alpha_{q})$, which is equal to zero only if $\alpha_{q}-\frac{1}{2}$ is an integer, which is impossible.
@@
Now assume, for the sake of contradiction, that $T(e_{0}h)$ is nonzero. Letting $S(z)$ be the polynomial such that $R(z)=S(z-\frac{n}{2})$ and using the evenness of $R$, we obtain
$T(a\rho(a))$ $=T(e_{0}h)(R(n)\bar{R}(0)+R(0)\bar{R}(n))\prod\limits_{q=0}^{n-1}(c_{0}q+\beta_{q})$
$\phantom{=\ }+\int_{i\mathbb{R}}R(-z)\bar{R}(z+n)\left(\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})\right)w_{0}(z)|\mathrm{d}z|$
$=T(e_{0}h)(S(\tfrac{n}{2})\bar{S}(-\tfrac{n}{2})+S(-\tfrac{n}{2})\bar{S}(\tfrac{n}{2}))\prod\limits_{q=0}^{n-1}(c_{0}q+\beta_{q})$
@@
$\phantom{=\ }+\int_{i\mathbb{R}}S(-z-\tfrac{n}{2})\bar{S}(z+\tfrac{n}{2})\left(\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})\right)w_{0}(z)|\mathrm{d}z|$
$=T(e_{0}h)(S(\tfrac{n}{2})\bar{S}(-\tfrac{n}{2})+S(-\tfrac{n}{2})\bar{S}(\tfrac{n}{2}))\prod\limits_{q=0}^{n-1}(c_{0}q+\beta_{q})$
$\phantom{=\ }+\int_{i\mathbb{R}+\frac{n}{2}}S(-z)\bar{S}(z)w(z)|\mathrm{d}z|,$
[…]
$T(a\rho(a))=\psi(S)+\int_{i\mathbb{R}}S(-z)\bar{S}(z)w(z)|\mathrm{d}z|,$
@@
where $\psi(S)$ is proportional to $S(\frac{n}{2})\bar{S}(-\frac{n}{2})+S(-\frac{n}{2})\bar{S}(\frac{n}{2})$. By suitable choice of $S$, we can force $\psi(S)$ to be negative; let $S_{0}$ be a polynomial with this property and satisfying that $S_{0}(z)=S_{0}(n-z)$, and let $U(z)=(z-\frac{n}{2})^{2}$, so that $\psi(S_{0}-UL)=\psi(S_{0})$ for any polynomial $L$ and $U(z)=U(n-z)$. By Lemma 1.5.1(1), there exists a sequence of polynomials $S_{i}$ such that $US_{i}$ tends to $S_{0}$ in the space $L^{2}(i\mathbb{R},|w(z)|+|w(n-z)|)$. It follows from the symmetry of $S_{0}$ and $U$ that $U(z)\cdot\frac{S_{i}(z)+S_{i}(n-z)}{2}$ tends to $S_{0}$ in the space $L^{2}(i\mathbb{R},|w(z)|)$. Therefore, the integral
$\int_{i\mathbb{R}}\left(S_{0}(-z)-U(-z)\cdot\frac{S_{i}(-z)+S_{i}(n+z)}{2}\right)\left(\bar{S}_{0}(z)-\bar{U}(z)\cdot\frac{\bar{S}_{i}(z)+\bar{S}_{i}(n-z)}{2}\right)w(z)|\mathrm{d}z|$
@@
approaches 0. On the other hand, $\psi(S_{0}(z)-U(z)\cdot\frac{S_{i}(z)+S_{i}(n-z)}{2})=\psi(S_{0})$ is negative. Therefore, choosing $S(z)=S_{0}(z)-U(z)\cdot\frac{S_{i}(z)+S_{i}(n-z)}{2}$ for large enough $i$ and noting that the resulting polynomial $R$ is even, we obtain $T(a\rho(a))<0$, a contradiction. ∎
@@
Let $a=(x^{n}+y^{n}+S(k))e$ for some even polynomial $S$. As in Proposition 1.5.3, $ea=a$, so $a$ is in $\mathcal{O}_{c}$, and we must have $T(a\rho(a))=0$. Expanding and using the evenness of $S$ gives
$T(a\rho(a))$ $=T\left((x^{n}+y^{n}+S(k))e(x^{n}+y^{n}+\bar{S}(k))e\right)$
$=T\left((x^{n}+y^{n}+S(k))(x^{n}+y^{n}+\bar{S}(k))e\right)$
$=T\left((x^{2n}+x^{n}y^{n}+y^{n}x^{n}+y^{2n})e\right)+T\left(S(k)\bar{S}(k)e\right)$
$\phantom{=\ }+T\left(x^{n}(S(k-n)+\bar{S}(k))e\right)+T\left(y^{n}(S(k+n)+\bar{S}(k))e\right).$
We evaluate these terms separately. By Proposition 1.5.3 and Section 1.5 with $R=1$, we have
$T\left((x^{2n}+x^{n}y^{n}+y^{n}x^{n}+y^{2n})e\right)$ $=\int_{i\mathbb{R}}\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})w_{0}(z)|\mathrm{d}z|.$
Splitting into $e_{0}$ and $e_{0}h$ components and using Proposition 1.5.3, we find that the second term is simply
$T\left(S(k)\bar{S}(k)e\right)$ $=\tfrac{1}{2}T\left(S(k)\bar{S}(k)e_{0}\right)+\tfrac{1}{2}T\left(S(k)\bar{S}(k)e_{0}h\right)$
$=\int_{i\mathbb{R}}S(z)\bar{S}(z)w_{0}(z)|\mathrm{d}z|.$
Again using Proposition 1.5.3 and applying Conditions (0) and (0) of Theorem 1.3.2, we find that the sum of the last two terms is equal to
$\tfrac{1}{2}T\left(x^{n}(S(k-n)+\bar{S}(k))e_{0}h\right)+\tfrac{1}{2}T\left(y^{n}(S(k+n)+\bar{S}(k))e_{0}h\right)$
$=\tfrac{1}{2}T\left(x^{m}y^{m}(S(k-m)+\bar{S}(k+m))e_{m}h\right)+(-1)^{m}\tfrac{1}{2}T\left(y^{m}x^{m}(S(k+m)+\bar{S}(k-m))e_{m}h\right),$
@@
$T(x^{m}y^{m}(S(k-m)+\bar{S}(k+m))e_{m}h)=T(e_{m}h)(S(-m)+\bar{S}(m))\prod\limits_{q=0}^{m-1}(c_{0}q+\beta_{m+q}).$
@@
$T(a\rho(a))$ $=T(e_{m}h)(S(-m)+\bar{S}(m))\prod\limits_{q=0}^{m-1}(c_{0}q+\beta_{m+q})$
$\phantom{=\ }+\int_{i\mathbb{R}}\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})w_{0}(z)|\mathrm{d}z|$
$\phantom{=\ }+\int_{i\mathbb{R}}S(-z)\bar{S}(z)w_{0}(z)|\mathrm{d}z|.$
Assume, for the sake of contradiction, that $T(e_{m}h)$ is nonzero. As in the proof of Proposition 1.5.3, the product $\prod_{q=0}^{m-1}(c_{0}q+\beta_{m+q})$ is nonzero by Lemma 1.2.5 and the condition on the $\alpha_{q}$, so for a suitable choice $S_{0}$ of even $S$, the expression
$T(e_{m}h)(S(-m)+\bar{S}(m))\prod\limits_{q=0}^{m-1}(c_{0}q+\beta_{m+q})+\int_{i\mathbb{R}}\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})w_{0}(z)|\mathrm{d}z|$
can be made negative. Taking $U(z)=(z-m)(z+m)$, there exists a sequence $S_{i}$ of polynomials such that $US_{i}$ approaches $S_{0}$ in $L^{2}(i\mathbb{R},|w_{0}(z)|)$ by Lemma 1.5.1(1), so $U(z)\cdot\frac{S_{i}(z)+S_{i}(-z)}{2}$ approaches $S_{0}$ in $L^{2}(i\mathbb{R},|w_{0}(z)|)$ because $w_{0}$ is even. Setting $S(z)=S_{0}(z)-U(z)\cdot\frac{S_{i}(z)+S_{i}(-z)}{2}$ for large enough $i$, it follows that $T(a\rho(a))<0$, a contradiction. ∎
@@
By this inclusion, any trace on $\mathcal{O}_{c}^{C_{n}}$ can be restricted to a trace on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$. Observing that the formula in Theorem 1.4.2 is identical to the analytic formula for type A traces obtained in (EKRS, 21, Proposition 3.1) with $z=\frac{k}{n}e_{0}$ and $u$ and $v$ scalar multiples of $\frac{x^{n}}{c_{0}n}e_{0}$ and $\frac{y^{n}}{c_{0}n}e_{0}$, respectively, such that the polynomial $P$ satisfying $uv=P(z+\frac{1}{2})$ obtained from Lemma 1.2.4 is monic, we see that the resulting restricted trace $T$ can be any trace satisfying Theorem 1.4.2 and the additional condition that $T(e_{q}h)=0$ for all $q$. This condition follows from Lemma 1.3.1 and Propositions 1.5.3 and 1.5.4, so any positive trace on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$ is the restriction of some trace on $\mathcal{O}_{c}^{C_{n}}$.

### v2

@@ 6 Positive traces
    If $H(z)$ is a continuous complex-valued function on $\mathbb{R}$ with finitely many zeros and at most polynomial growth at infinity, then the set $\{H(z)S(z)\mid S(z)\in\mathbb{C}[z]\}$ is dense in the space $L^{p}(\mathbb{R},w)$.
@@
    The closure of the set $\{S(z)\bar{S}(z)\mid S(z)\in\mathbb{C}[z]\}$ in $L^{p}(\mathbb{R},w)$ is the subset of almost everywhere nonnegative functions.
We begin by showing that the linear functional $\Phi_{0}$ of Theorem 5.3 must be identically zero.
Proposition 6.2.
Any positive trace $T$ on $\mathcal{O}_{c}$ of the form given by Theorem 5.3 satisfies $\Phi_{0}(R)=0$} for all polynomials $R$.
@@
Since $T$ is a positive trace, we must have $T(a\rho(a))>0$ for all $a$ in $\mathcal{O}_{c}$. In particular, we must have $T\big(R(k)\bar{R}(k)e\big)>0$ for all even polynomials $R$. By condition (0) of Theorem 4.2, this trace value is equal to $\frac{1}{2}T\big(R(k)\bar{R}(k)e_{0}\big)+\frac{1}{2}R(0)\bar{R}(0)T(e_{0}h)$.
Suppose, for the sake of contradiction, that $\Phi_{0}$ is not identically zero. Then there exists a polynomial $S_{0}$ such that $S_{0}(0)=0$ and $\Phi_{0}(S_{0})$ is negative. By Remark 5.4, we may take $S_{0}$ to be even. Also, let
@@
where the product is over singularities $z_{0}$ of $w_{0}(z)$ between $\rm i$$\mathbb{R}$ and the path $C$, so that $H(0)=0$ and $H\bigl(\alpha_{i}-i-\frac{1}{2}\bigr)=0$ for all $i$. Thus for any polynomial $U(z)$, we have $\Phi_{0}\bigl((S_{0}-HU)\bigl(\bar{S}_{0}-\bar{H}\bar{U}\bigr)\bigr)=\Phi_{0}(S_{0})$ and $S_{0}(0)-H(0)U(0)=0$. Also, by Remark 5.4, the polynomial $H(z)$ is even.
By part (1) of Lemma 6.1, there exists a sequence of polynomials $S_{i}$ such that $HS_{i}$ tends to $S_{0}$ in the space $L^{2}({\rm i}\mathbb{R},|w_{0}|)$. Because $w_{0}(z)$ is even, the sequence $H(z)\cdot\frac{S_{i}(z)+S_{i}(-z)}{2}$ also tends to $S_{0}$ in $L^{2}({\rm i}\mathbb{R},|w_{0}|)$. Thus the integral
$\int_{{\rm i}\mathbb{R}}\left(S_{0}(z)-H(z)\cdot\frac{S_{i}(z)+S_{i}(-z)}{2}\right)\left(\bar{S}_{0}(z)-\bar{H}(z)\cdot\frac{\bar{S}_{i}(z)+\bar{S}_{i}(-z)}{2}\right)w_{0}(z)\mathrm{d}z$
@@
$T\bigl(R(k)\bar{R}(k)e\bigr)={}$ $\tfrac{1}{2}T\bigl(R(k)\bar{R}(k)e_{0}\bigr)+\tfrac{1}{2}R(0)\bar{R}(0)T(e_{0}h)=\tfrac{1}{2}T\bigl(R(k)\bar{R}(k)e_{0}\bigr)$
$={}$ $\int_{{\rm i}\mathbb{R}}R(z)\bar{R}(z)w_{0}(z)\mathrm{d}z+\Phi_{0}(R),$
@@
$T(a\rho(a))={}$ $T\bigl(R(k)(x^{n}+y^{n})e\bar{R}(k)(x^{n}+y^{n})e\bigr)$
$={}$ $T\bigl(R(k)(x^{n}+y^{n})\bar{R}(k)(x^{n}+y^{n})e\bigr)$
$={}$ $T\bigl(R(k)(x^{n}+y^{n})\bigl(x^{n}\bar{R}(k-n)+y^{n}\bar{R}(k+n)\bigr)e\bigr),$
and distributing and using the commutativity relations with $k$ listed in Lemma 3.1, we obtain
$T(a\rho(a))={}$ $T\bigl(R(k)\bigl(x^{2n}\bar{R}(k-n)+x^{n}y^{n}\bar{R}(k+n)+y^{2n}\bar{R}(k+n)+y^{n}x^{n}\bar{R}(k-n)\bigr)e\bigr)$
$={}$ $T\bigl(x^{2n}R(k-2n)\bar{R}(k-n)e\bigr)+T\bigl(x^{n}y^{n}R(k)\bar{R}(k+n)e\bigr)$
$+T\bigl(y^{2n}R(k+2n)\bar{R}(k+n)e\bigr)+T\bigl(y^{n}x^{n}R(k)\bar{R}(k-n)e\bigr).$
@@
$T\bigl(a\rho(a)\bigr)={}$ $\tfrac{1}{2}T\bigl(x^{2n}R(k-2n)\bar{R}(k-n)e_{0}h\bigr)+\tfrac{1}{2}T\bigl(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}h\bigr)$
$+\tfrac{1}{2}T\bigl(y^{2n}R(k+2n)\bar{R}(k+n)e_{0}h\bigr)+\tfrac{1}{2}T\bigl(y^{n}x^{n}R(k)\bar{R}(k-n)e_{0}h\bigr)$
$+\tfrac{1}{2}T\bigl(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}\bigr)+\tfrac{1}{2}T\bigl(y^{n}x^{n}R(k)\bar{R}(k-n)e_{0}\bigr).$
Using conditions (0) and (0) of Theorem 4.2 and rearranging terms, we obtain
$T(a\rho(a))={}$ $\tfrac{1}{2}T\bigl(x^{n}y^{n}R(k-n)\bar{R}(k)e_{0}h\bigr)+\tfrac{1}{2}T\bigl(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}h\bigr)$
$+\tfrac{1}{2}T\bigl(y^{n}x^{n}R(k+n)\bar{R}(k)e_{0}h\bigr)+\tfrac{1}{2}T\bigl(y^{n}x^{n}R(k)\bar{R}(k-n)e_{0}h\bigr)$
$+\tfrac{1}{2}T\bigl(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}\bigr)+\tfrac{1}{2}T\bigl(y^{n}x^{n}R(k)\bar{R}(k-n)e_{0}\bigr).$
@@
$T\bigl(a\rho(a)\bigr)={}$ $T\bigl(x^{n}y^{n}e_{0}h\bigr)\bigl(R(n)\bar{R}(0)+R(0)\bar{R}(n)\bigr)+T\bigl(x^{n}y^{n}R(k)\bar{R}(k+n)e_{0}\bigr)$
$={}$ $T(e_{0}h)\bigl(R(n)\bar{R}(0)+R(0)\bar{R}(n)\bigr)\prod\limits_{q=0}^{n-1}(c_{0}q+\beta_{q})$
$+\int_{{\rm i}\mathbb{R}}R(z)\bar{R}(z+n)\left(\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})\right)w_{0}(z)|\mathrm{d}z|.$ (6.1)
@@
The product $\prod_{q=0}^{n-1}(c_{0}q+\beta_{q})$ in the first term is nonzero by assumption; indeed, by Lemma 3.5, $c_{0}q+\beta_{q}=c_{0}\bigl(q+\frac{1}{2}-\alpha_{q}\bigr)$, which is equal to zero only if $\alpha_{q}-\frac{1}{2}$ is an integer, which is impossible.
@@
Now assume, for the sake of contradiction, that $T(e_{0}h)$ is nonzero. Letting $S(z)$ be the polynomial such that $R(z)=S\bigl(z-\frac{n}{2}\bigr)$ and using the evenness of $R$, we obtain
$T(a\rho(a))={}$ $T(e_{0}h)\bigl(R(n)\bar{R}(0)+R(0)\bar{R}(n)\bigr)\prod\limits_{q=0}^{n-1}(c_{0}q+\beta_{q})$
$+\int_{{\rm i}\mathbb{R}}R(-z)\bar{R}(z+n)\left(\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})\right)w_{0}(z)|\mathrm{d}z|$
$={}$ $T(e_{0}h)\bigl(S\bigl(\tfrac{n}{2}\bigr)\bar{S}\bigl(-\tfrac{n}{2}\bigr)+S\bigl(-\tfrac{n}{2}\bigr)\bar{S}\bigl(\tfrac{n}{2}\bigr)\bigr)\prod\limits_{q=0}^{n-1}(c_{0}q+\beta_{q})$
@@
$+\int_{{\rm i}\mathbb{R}}S\bigl(-z-\tfrac{n}{2}\bigr)\bar{S}\bigl(z+\tfrac{n}{2}\bigr)\left(\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})\right)w_{0}(z)|\mathrm{d}z|$
$={}$ $T(e_{0}h)\bigl(S\bigl(\tfrac{n}{2}\bigr)\bar{S}\bigl(-\tfrac{n}{2}\bigr)+S\bigl(-\tfrac{n}{2}\bigr)\bar{S}\bigl(\tfrac{n}{2}\bigr)\bigr)\prod\limits_{q=0}^{n-1}(c_{0}q+\beta_{q})$
$+\int_{{\rm i}\mathbb{R}+\frac{n}{2}}S(-z)\bar{S}(z)w(z)|\mathrm{d}z|,$
where $w$ is a function with $|w(z)|$ satisfying the conditions of Lemma 6.1 (1). Because $\beta_{q}=\frac{c_{0}}{2}-c_{0}\alpha_{q}$ by Lemma 3.5, the weight function $w(z)$ is holomorphic between $\rm i$$\mathbb{R}$ and ${\rm i}\mathbb{R}+n$. Moving the contour of integration, we obtain
$T(a\rho(a))=\psi(S)+\int_{{\rm i}\mathbb{R}}S(-z)\bar{S}(z)w(z)|\mathrm{d}z|,$
@@
where $\psi(S)$ is proportional to $S(\frac{n}{2})\bar{S}\bigl(-\frac{n}{2}\bigr)+S\bigl(-\frac{n}{2}\bigr)\bar{S}\bigl(\frac{n}{2}\bigr)$. By suitable choice of $S$, we can force $\psi(S)$ to be negative; let $S_{0}$ be a polynomial with this property and satisfying that $S_{0}(z)=S_{0}(n-z)$, and let $U(z)=\bigl(z-\frac{n}{2}\bigr)^{2}$, so that $\psi(S_{0}-UL)=\psi(S_{0})$ for any polynomial $L$ and U(z)=U(nz)U(z)=U(n-z). By Lemma 6.1 (1), there exists a sequence of polynomials $S_{i}$ such that $US_{i}$ tends to $S_{0}$ in the space $L^{2}({\rm i}\mathbb{R},|w(z)|+|w(n-z)|)$. It follows from the symmetry of $S_{0}$ and $U$ that $U(z)\cdot\frac{S_{i}(z)+S_{i}(n-z)}{2}$ tends to $S_{0}$ in the space $L^{2}({\rm i}\mathbb{R},|w(z)|)$. Therefore, the integral
$\int_{{\rm i}\mathbb{R}}\left(S_{0}(-z)-U(-z)\cdot\frac{S_{i}(-z)+S_{i}(n+z)}{2}\right)$
$\qquad\times\left(\bar{S}_{0}(z)-\bar{U}(z)\cdot\frac{\bar{S}_{i}(z)+\bar{S}_{i}(n-z)}{2}\right)w(z)|\mathrm{d}z|$
@@
approaches 0. On the other hand, $\psi\bigl(S_{0}(z)-U(z)\cdot\frac{S_{i}(z)+S_{i}(n-z)}{2}\bigr)=\psi(S_{0})$ is negative. Therefore, choosing $S(z)=S_{0}(z)-U(z)\cdot\frac{S_{i}(z)+S_{i}(n-z)}{2}$ for large enough $i$ and noting that the resulting polynomial $R$ is even, we obtain $T(a\rho(a))<0$, a contradiction. ∎
@@
Let $a=(x^{n}+y^{n}+S(k))e$ for some even polynomial $S$. As in Proposition 6.3, $ea=a$, so $a$ is in $\mathcal{O}_{c}$, and we must have $T(a\rho(a))=0$. Expanding and using the evenness of $S$ gives
$T(a\rho(a))={}$ $T\bigl((x^{n}+y^{n}+S(k))e\bigl(x^{n}+y^{n}+\bar{S}(k)\bigr)e\bigr)$
$={}$ $T\bigl((x^{n}+y^{n}+S(k))(x^{n}+y^{n}+\bar{S}(k))e\bigr)$
$={}$ $T\bigl(\bigl(x^{2n}+x^{n}y^{n}+y^{n}x^{n}+y^{2n}\bigr)e\bigr)+T\bigl(S(k)\bar{S}(k)e\bigr)$
$+T\bigl(x^{n}\bigl(S(k-n)+\bar{S}(k)\bigr)e\bigr)+T\bigl(y^{n}(S(k+n)+\bar{S}(k))e\bigr).$
We evaluate these terms separately. By Proposition 6.3 and (6) with $R=1$, we have
$T\bigl((x^{2n}+x^{n}y^{n}+y^{n}x^{n}+y^{2n})e\bigr)={}$ $\int_{{\rm i}\mathbb{R}}\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})w_{0}(z)|\mathrm{d}z|.$
Splitting into $e_{0}$ and $e_{0}h$ components and using Proposition 6.3, we find that the second term is simply
$T\bigl(S(k)\bar{S}(k)e\bigr)={}$ $\tfrac{1}{2}T\bigl(S(k)\bar{S}(k)e_{0}\bigr)+\tfrac{1}{2}T\bigl(S(k)\bar{S}(k)e_{0}h\bigr)=\int_{{\rm i}\mathbb{R}}S(z)\bar{S}(z)w_{0}(z)|\mathrm{d}z|.$
Again using Proposition 6.3 and applying conditions (0) and (0) of Theorem 4.2, we find that the sum of the last two terms is equal to
$\tfrac{1}{2}T\bigl(x^{n}(S(k-n)+\bar{S}(k))e_{0}h\bigr)+\tfrac{1}{2}T\bigl(y^{n}(S(k+n)+\bar{S}(k))e_{0}h\bigr)$
$\qquad=\tfrac{1}{2}T\bigl(x^{m}y^{m}(S(k-m)+\bar{S}(k+m))e_{m}h\bigr)$
$\phantom{\qquad=}{}+(-1)^{m}\tfrac{1}{2}T\bigl(y^{m}x^{m}(S(k+m)+\bar{S}(k-m))e_{m}h\bigr),$
@@
$T\bigl(x^{m}y^{m}\bigl(S(k-m)+\bar{S}(k+m)\bigr)e_{m}h\bigr)=T(e_{m}h)\bigl(S(-m)+\bar{S}(m)\bigr)\prod\limits_{q=0}^{m-1}(c_{0}q+\beta_{m+q}).$
@@
$T(a\rho(a))={}$ $T(e_{m}h)\bigl(S(-m)+\bar{S}(m)\bigr)\prod\limits_{q=0}^{m-1}(c_{0}q+\beta_{m+q})+\int_{{\rm i}\mathbb{R}}\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})w_{0}(z)|\mathrm{d}z|$
$+\int_{{\rm i}\mathbb{R}}S(-z)\bar{S}(z)w_{0}(z)|\mathrm{d}z|.$
Assume, for the sake of contradiction, that $T(e_{m}h)$ is nonzero. As in the proof of Proposition 6.3, the product $\prod_{q=0}^{m-1}(c_{0}q+\beta_{m+q})$ is nonzero by Lemma 3.5 and the condition on the $\alpha_{q}$, so for a suitable choice $S_{0}$ of even $S$, the expression
$T(e_{m}h)\bigl(S(-m)+\bar{S}(m)\bigr)\prod\limits_{q=0}^{m-1}(c_{0}q+\beta_{m+q})+\int_{{\rm i}\mathbb{R}}\prod\limits_{q=0}^{n-1}(c_{0}(z+q)+\beta_{q})w_{0}(z)|\mathrm{d}z|$
can be made negative. Taking U(z)=(zm)(z+m)U(z)=(z-m)(z+m), there exists a sequence $S_{i}$ of polynomials such that $US_{i}$ approaches $S_{0}$ in $L^{2}({\rm i}\mathbb{R},|w_{0}(z)|)$ by Lemma 6.1 (1), so $U(z)\cdot\frac{S_{i}(z)+S_{i}(-z)}{2}$ approaches $S_{0}$ in $L^{2}({\rm i}\mathbb{R},|w_{0}(z)|)$ because $w_{0}$ is even. Setting
$S(z)=S_{0}(z)-U(z)\cdot\frac{S_{i}(z)+S_{i}(-z)}{2}$
for large enough $i$, it follows that $T(a\rho(a))<0$, a contradiction. ∎
@@
By this inclusion, any trace on $\mathcal{O}_{c}^{C_{n}}$ can be restricted to a trace on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$. Observing that the formula in Theorem 5.2 is identical to the analytic formula for type A traces obtained in [11, Proposition 3.1] with $z=\frac{k}{n}e_{0}$ and $u$ and $v$ scalar multiples of $\frac{x^{n}}{c_{0}n}e_{0}$ and $\frac{y^{n}}{c_{0}n}e_{0}$, respectively, such that the polynomial $P$ satisfying $uv=P\bigl(z+\frac{1}{2}\bigr)$ obtained from Lemma 3.4 is monic, we see that the resulting restricted trace $T$ can be any trace satisfying Theorem 5.2 and the additional condition that $T(e_{q}h)=0$ for all $q$. This condition follows from Lemma 4.1 and Propositions 6.3 and 6.4, so any positive trace on $\mathcal{O}_{c}^{\mathbb{D}_{n}}$ is the restriction of some trace on $\mathcal{O}_{c}^{C_{n}}$.